OpenStudy (anonymous):

simplify: (square root)100/81

OpenStudy (akshay_budhkar):

ok its simple, do you know the square root of 100 and 81?

OpenStudy (anonymous):



OpenStudy (jamesj):

\[ \sqrt{\frac{100}{81}} = \frac{\sqrt{100}}{\sqrt{81}} = \frac{ ...}{ ... } \]

OpenStudy (anonymous):

yes 10/9

OpenStudy (anonymous):

it can be written as  the square root of 100 is 10, and the square root of 81 is 9, therefore it would be 10/9